an investor invested a total of $2,300 in two mutual funds. One fund earned a 8% profit while the other earned a 5% profit. If the investor's total profit was $175, how much was invested in each mutual fund?
Answer by mananth(16946)

Part I 8.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
2300
Interest----- 175

Part I 8.00% per annum ---x *
Part II 5.00% per annum ---y
Total investment
x + 1 y= 2300 -------------1
Interest on both investments
8.00% x + 5.00% y= 175
Multiply by 100
8 x + 5 y= 17500.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -18400.00
Add this to (2)
0 x -3 y= -900
divide by -3
y = 300 - x)
Part I 8.00% $ 2000
Part II 5.00% $ 300

CHECK
2000 --------- 8.00% ------- 160.00
300 ------------- 5.00% ------- 15.00
Total -------------------- 175.00
11.6
m.ananth@hotmail.ca 1520